OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation x + 2y = 4. A: x-intercept: (0, 2), y-intercept: (4, 0) B: x-intercept: (2, 0), y-intercept: (4, 0) C: x-intercept: (0, 4), y-intercept: (2, 0) D: x-intercept: (4, 0), y-intercept: (0, 2)

OpenStudy (anonymous):

To solve for the x intercept, set y=0 and solve for x. To solve for the y intercept, set x=0 and solve for y. What do you get?

OpenStudy (anonymous):

I Didn't get anything because i really don't understand how to solve the problem

OpenStudy (anonymous):

Ok, so lets solve for the x intercept. So we'll plug in 0 for the x in the equation: x+2y=4 0+2y=4 2y=4 y=2 So your x intercept is 2. Now do the same thing but plug in 0 for y instead of x.

OpenStudy (anonymous):

im not graphing anything, im just choosing the correct x-intercept and y-intercept of the equation x + 2y = 4.

OpenStudy (anonymous):

I know, I didn't graph anything. Do you see how I found the y intercept?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

To find the y intercept, we replace "x" in the equation with "0" x+2y=4 0+2y=4 (we just replaced x with 0 and now we solve for y) 2y=4 (divide both sides by 2) y=2

OpenStudy (anonymous):

OHHHH Ok so the answer is D?

OpenStudy (anonymous):

Right, exactly